OpenStudy (anonymous):

b + 2 over 3 - b - 2 over 7 = 2 Solve the equation. b + 2 over 3 - b - 2 over 7 = 2 Solve the equation. @Mathematics

OpenStudy (anonymous):

Find the common denominator. \[\left(\begin{matrix}7b+14\\21 \end{matrix}\right)\] - \[\left(\begin{matrix}9-3b-2 \\ 21\end{matrix}\right)\] Combine like terms \[\left(\begin{matrix}7b-3b+14-9-2 \\ 21\end{matrix}\right)\] \[\left(\begin{matrix}4b+3 \\ 21\end{matrix}\right)=2\] b=4/21

jhonyy9 (jhonyy9):

7b+14-3b-6=42 4b=34 2b=17 b=17/2